OpenStudy (blu_seaa):

What is √252 in simplified radical form? A. 21√12 B. 12√21 C. 7√6 D. 6√7 I think it's D.

OpenStudy (luckyness):

You are correct! The answer is D. Let me type out the reason of why~

OpenStudy (luckyness):

Factor 252 into 36*7 Break up the square roots using the identity \[\sqrt{x*y}=\sqrt{x}*\sqrt{y}\] Take the square root of the perfect square 36 to get 6 And you get \[6\sqrt{7}\] ~Luckyness

OpenStudy (blu_seaa):

Thank you. c: